# X-ray free electron laser SPI single-particle diffractive imaging

### v1

@@ 1 Introduction
Generative diffusion models have strongly impacted the field of machine learning in the past few years, with widespread applications from unconditional and conditional image and audio generation [35, 65, 61, 45] to data-driven approaches for solving inverse problems [13, 40, 75, 51]. In this work, we use score-based generative models, a subclass of diffusion models formulated via a continuous-time diffusion process, as data-driven priors, which have recently begun to be employed in real imaging problems, for example in a Plug-and-Play (PnP)-based method for (non-blind) ptychography [18]. We compare these data-driven priors against the use of either no prior information or a simple model-based total variation (TV) prior. The comparison is carried out using two algorithmic frameworks for variational inference [24, 51] suited for model-based or diffusion-based priors, as well as simpler optimization-based procedures.
@@ 2.2 Ptychography
Ptychography can also be interpreted as a specific type of coded illumination or coded aperture method, making links to other imaging disciplines [19], or as a variant of the short-time Fourier transform phase retrieval problem, with links to signal and audio processing [30, 63, 29].
@@ 3.1.3 RED-Diff
A simple variational distribution is a Gaussian with mean $\mu\in\mathbb{R}^{d}$ and isotropic covariance with a scalar $\sigma>0$, i.e., $q_{\phi}=\mathcal{N}(\mu,\sigma^{2}I)$, with $\phi=(\mu,\sigma)$. In [51], the authors prove that the VI objective (3.9) can then be written as
@@ 3.1.4 Variational inference with principled score-based priors
In two other works [25, 24], the VI loss is optimized without reweighting in time. Evaluating the objective (or its gradients) in (3.9) requires evaluating the prior log-density $\log{\hat{p}_{\theta}}(x)$ for unseen data $x$. The score model $s_{\theta}(x_{t},t)$ is typically unstable around $t\approx 0$, but as previously shown by Song et al. [65], log-density values can instead be obtained by solving the initial value problem for the forward probability flow ordinary differential equation (ODE)
$\frac{\mathrm{d}x_{t}}{\mathrm{d}t}=f(x_{t},t)-\frac{1}{2}g(t)^{2}\nabla_{x}\log p_{t}(x_{t},t),\hskip 18.49988ptx_{0}=x.$ (3.11)
@@
This generates the same dynamics of the distribution $p_{t}$ as (3.5), since $p_{t}$ solves both the continuity equation for (3.11) as well as the Fokker–Planck equation of (3.5). Replacing the score in (3.11) by the score model $s_{\theta}$ allows solving this ODE, giving an approximation $\log{\hat{p}_{\theta}}^{\mathrm{ODE}}\approx\log{\hat{p}_{\theta}}$. The authors of [25] thus proposed to use $\log{\hat{p}_{\theta}}^{\mathrm{ODE}}$ to evaluate the objective (3.9). As the numerical results of [25] showed, the resulting VI approach has a high computational cost, but generates very accurate approximations of the true posterior.
A follow-up paper [24] resolves the costly evaluation of $\log{\hat{p}_{\theta}}^{\mathrm{ODE}}$ by instead employing the evidence lower bound (ELBO) surrogate $b_{\theta}^{\mathrm{SDE}}\leq\log{\hat{p}_{\theta}}^{\mathrm{ODE}}$; the exact definition of $b_{\theta}^{\mathrm{SDE}}$ can be found in [24, p. 5]. For the readers convenience, it is repeated in Appendix B. After this modification, one instead solves
@@
$\tilde{\phi}\coloneqq\operatorname*{arg\,min}_{\phi}\left\{\mathbb{E}_{x\sim q_{\phi}}\left[-\log p(y\,|\,x)-b_{\theta}^{\mathrm{SDE}}(x)\right]-\mathcal{H}(q_{\phi})\right\}.$ (3.12)
@@
where the ELBO term $b_{\theta}^{\mathrm{SDE}}$ can be estimated efficiently using Monte Carlo integration. Heuristically, the ELBO term induces a surrogate score-based prior (SSP) with density ${\hat{p}_{\theta}}^{\mathrm{surr}}\propto\exp(b_{\theta}^{\mathrm{SDE}})$. Empirically, it was shown that this surrogate prior is very effective, in the sense that for small dimensional examples $\tilde{\phi}\approx\phi^{\ast}$, i.e., the fitted distribution shows very good agreement with the variational approximation to the posterior. The reduced computational time of (3.12) allows to lift the problem dimension to realistic imaging sizes.
@@
We now describe how we can extend this method to the case of blind inverse problems (3.4), using a standard variational Bayes approach for the joint posterior of image $x$ and latent parameter $\mathbf{r}$. While we are specifically interested in solving the position-blind ptychography problem (4.2), the method we develop can be used for general blind imaging inverse problems.
@@ 3.2.1 Generalized VI Approach for Blind Problems
Using (3.4), we generalize the VI approach to the joint posterior $p(x,\mathbf{r}\,|\,y)$ in the blind case. Under the non-restrictive assumption of independent priors $\log p(x,\mathbf{r})=\log{\hat{p}_{\theta}}(x)+\log p(\mathbf{r})$, the derivation of the VI objective (3.9) easily generalizes to
@@
Note that the variational distribution $q_{\phi}$ is now the joint distribution of image and parameters. Depending on the forward model and dimensionality, the joint optimization problem can exhibit strong non-convexity and be computationally demanding. The problem can be provided with more structure by a standard mean-field assumption on the variational Bayesian posterior
@@
where $\phi=(\chi,\rho)$, with $\chi$ encoding the image posterior and $\rho$ the latent parameter posterior. Note that this comes at the cost of less accurately representing the correlation of uncertainties in image and parameters in $q_{\phi}$, but is necessary to make the optimization feasible at scale. We note that a detailed analysis of this approximation is an important open question for future research, but is outside of the scope of this work. By inspecting the optimality conditions of $\min_{\phi}L(\phi)=\min_{\chi,\rho}L(\chi,\rho)$ with respect to $\chi$ and $\rho$ separately, under the mean-field assumption, the posterior marginals obey the optimality conditions [7]
@@
$\rho_{k+1}$ $=\operatorname*{arg\,min}_{\chi}\mathcal{L}(\chi_{k+1},\rho)=\operatorname*{arg\,min}_{\rho}\left\{\mathbb{E}_{x\sim q_{\chi_{k+1}}}\mathbb{E}_{\mathbf{r}\sim q_{\rho}}\left[-\log p(y\,|\,x,\mathbf{r})-\log p(\mathbf{r})\right]-\mathcal{H}(q_{\rho})\right\}.$ (3.15b)
@@ 3.2.2 Blind RED-Diff
The RED-Diff method has been modified for tackling blind imaging inverse problems on the example of MRI with unknown off-resonance field map [1]. The authors there employed the same mean-field assumption (3.14) in order to separate optimization steps for the reconstructed image and parameters. As before, the objective for the image posterior can be rewritten using [51, Prop. 1] to arrive at a loss similar to (3.10), but with the likelihood conditioned on the current parameter estimate, i.e., (3.15a) becomes
@@
with a suitable weight $\tilde{\omega}(t)$; see [51, Prop. 2] for details. For the parameter update step (3.15b), the authors of [1] showed that assuming $p(\mathbf{r})$ and $q_{\rho}(\mathbf{r})$ are both Laplace distributions allows to obtain closed form representations of the relevant terms $p(\mathbf{r})$ and $\mathcal{H}(q_{\rho})$. Upon replacing $q_{\rho}$ with a point mass on a single $\mathbf{r}$, the entropy is, however, dropped again in [1] and the parameter update becomes
@@
We line out the resulting algorithm with our notation in Algorithm 2.
@@ 3.2.3 Blind surrogate score-based prior (SSP) method
$\chi_{k+1}=\operatorname*{arg\,min}_{\chi}\mathcal{L}(\chi,\rho_{k})=\operatorname*{arg\,min}_{\chi}\left\{\mathbb{E}_{x\sim q_{\chi}}\mathbb{E}_{\mathbf{r}\sim q_{\rho_{k}}}\left[-\log p(y\,|\,x,\mathbf{r})(x)-b_{\theta}^{\mathrm{SDE}}(x)\right]-\mathcal{H}(q_{\chi})\right\}$ (3.19)
For medium- to large-scale imaging problems, $q_{\chi}$ can be a Gaussian distribution with diagonal covariance – more complex models were computationally prohibitive in our setup; see also [24] for scaling experiments. We solve (3.15) using inner loops of stochastic gradient descent, where the partial derivatives $\nabla_{\chi}\mathcal{L}$ and $\nabla_{\rho}\mathcal{L}$ are approximated using Monte Carlo estimators of the expectations and the typical reparametrization trick [41] whenever $q_{\chi},q_{\rho}$ are Gaussian. This sampling-based approach admits the use of a batch size $B\geq 1$ for each gradient evaluation, which can reduce the variance of estimated gradients. The surrogate prior term $b_{\theta}^{\mathrm{SDE}}(x)$ can be implemented using a single forward-pass through $s_{\theta}$; see [24, sec. 4.2]. This leads to the variational Bayes approach summarized in Algorithm 1. For a clearer presentation, we formulate the updates as single gradient descent steps. However, in the practical implementation, lines 5 and 10 in Algorithm 1 are replaced by the Adam optimization scheme [42].

### v2

@@ 1 Introduction
Generative diffusion models have strongly impacted the field of machine learning in the past few years, with widespread applications from unconditional and conditional image and audio generation [57, 98, 89, 71] to data-driven approaches for solving inverse problems [25, 63, 110, 77]. Multiple prior works have also investigated the use of (both diffusion-based and not diffusion-based) generative models for ptychography specifically [93, 94, 92, 68]. In this work, we use score-based generative models, a subclass of diffusion models formulated via a continuous-time diffusion process, as data-driven priors, which have recently begun to be employed in real imaging problems, for example in a Plug-and-Play (PnP)-based method for (non-blind) ptychography [30]. We compare these data-driven priors against the use of either no prior information or a simple model-based total variation (TV) prior. The comparison is carried out using two algorithmic frameworks for variational inference [38, 77] suited for model-based or diffusion-based priors, as well as simpler optimization-based procedures. In both classes of methods, we make use of auto-differentiation, which has been successfully used in multiple works on ptychography [69, 91, 61, 47, 81].
Our key contributions in this work are as follows: (1) we introduce and discuss the novel and challenging blind inverse problem of position-blind ptychography; (2) we derive an extension of the modern generative surrogate score-based prior (SSP) variational reconstruction method [38] to the parameter-blind setting; (3) we run extensive numerical simulations, employing and comparing classic model-based and modern data-driven reconstruction methods in multiple variants of the position-blind ptychography problem; (4) we offer empirical insights into the inherent difficulties associated with position-blind ptychography, as well as possible remedies inspired by classic optimization ideas and recent phaseless imaging literature; (5) we make our code available to the research community at https://github.com/sp-uhh/PositionBlindPtycho.
@@ 2.2 Ptychography
Ptychography can also be interpreted as a specific type of coded illumination or coded aperture method, making links to other imaging disciplines [32, 2], or as a variant of the short-time Fourier transform phase retrieval problem, with links to signal and audio processing [46, 95, 45].
@@
Remark 2.1.
The ill-posedness of the ptychographic phase retrieval problem has been studied in multiple works, see [50, 22, 36, 42, 51]. For example, [50] proves that the solution is unique (up to a global phase factor) assuming the short-time Fourier transform of the probe function does not vanish, and gives sufficient conditions on the probe function to verify this. Results of this type, however, rely on knowledge of the measurement positions. In particular, well-posedness can be ensured given that there is enough overlap between measurements at different positions. In the position-blind setting, a similar assumptions seems problematic due to the stochastic nature of the measurement process and the scan positions. An open problem is to identify a condition that holds with high-probability under which uniqueness can be guaranteed. We later empirically support the high difficulty of the position-blind scenario by showing the high amount of non-convexity of the position recovery loss landscape, even in the idealized setting where the target object is already known, see Fig. 5.
@@ 2.3 Single-Particle Imaging and Ptychography
2.4 Position-blind ptychography
In this work, we will treat a simplified 2D far-field real-space position-blind ptychography problem. For clarity, we preface this section with the following simplifying assumptions we make for computational viability in our simulated setting:
Assumption 2.1.
Assumption 2.2.
Assumption 2.3.
Assumption 2.4.
Assumption 2.5.
These assumptions together lead to a simplified toy problem variant of the full 3-D problem, as one would have to treat many additional unknowns (e.g., the probe function, sample rotations) in a realistic SPI setting. However, due to the difficulty that our simplified problem setup already exhibits, and due to the need for parameter tuning in more complex settings, we leave relaxations of the above assumptions for future work. Nonetheless, we note that the reconstruction methods we present herein can be extended to this end, and provide concrete ideas for such future extensions in Section 6.1.
We now turn our attention to the concrete forward model we choose. Let $x\in\mathbb{C}^{H\times W}$ be an image of the complex object transmission function which we want to recover,
$y_{k}$ $=\left|\mathcal{A}(r_{k},x)\right|^{2}+\varepsilon_{k},$ (2.2)
where $\varepsilon_{k}$ is the measurement noise of the $k$-th measurement, $\mathcal{F}$ models the wavefront propagation to the detector and $\operatorname{DFT}$ denotes the discrete Fourier transform, the probe array $p\in\mathbb{C}^{H_{p}\times W_{p}}$ models the complex-valued wavefield of the probe in the plane of the object and is applied through element-wise multiplication $\odot$, and $S(r_{k},\cdot)$ is a shift operator depending on a two-dimensional shift $r_{k}=(\Delta h_{k},\Delta w_{k})$ which we refer to as the scan position in the following. The Crop operator crops its input to the same array size as the probe $p$ before element-wise multiplication. The Pad operator pads its input with one full probe array size of empty space entries; see Section 4.1.1.
@@ 3.1.3 RED-Diff
A simple variational distribution is a Gaussian with mean $\mu\in\mathbb{R}^{d}$ and isotropic covariance with a scalar $\sigma>0$, i.e., $q_{\phi}=\mathcal{N}(\mu,\sigma^{2}I)$, with $\phi=(\mu,\sigma)$. In [77], the authors prove that the VI objective (3.1.2) can then be written as
@@ 3.1.4 Variational inference with principled score-based priors
In two other works [39, 38], the VI loss is optimized without reweighting in time. Evaluating the objective (or its gradients) in (3.1.2) requires evaluating the prior log-density $\log{\hat{p}_{\theta}}(x)$ for unseen data $x$. The score model $s_{\theta}(x_{t},t)$ is typically unstable around $t\approx 0$, but as previously shown by Song et al. [98], log-density values can instead be obtained by solving the initial value problem for the forward probability flow ordinary differential equation (ODE)
$\frac{\mathrm{d}x_{t}}{\mathrm{d}t}=\underbrace{f(x_{t},t)-\frac{1}{2}g(t)^{2}\nabla_{x}\log p_{t}(x_{t},t)}_{=:v(x_{t},t)},\hskip 18.49988ptx_{0}=x.$ (3.11)
This generates the same dynamics of the distribution $p_{t}$ as (3.5), since the continuity equation of (3.11) coincides with the Fokker–Planck equation of (3.5). For velocity fields $v\in C^{2}$, [104, Cor. 10.4] proves the instantaneous change of variables formula
$\frac{\mathrm{d}}{\mathrm{d}t}p_{t}=-p_{t}(x_{t})\nabla\cdot v(x_{t},t)\qquad\Leftrightarrow\qquad\frac{\mathrm{d}}{\mathrm{d}t}\log p_{t}=-\nabla\cdot v(x_{t},t)$
see also [23]. We can thus write $\log p_{0}$ as the solution of an initial-value problem:
$\log p_{0}(x_{0})=\log p_{T}(x_{T})+\int_{0}^{T}\nabla\cdot v(x_{t},t)\mathrm{d}t$
@@
Replacing the score in (3.11) by a score model $s_{\theta}$ induces a learned velocity field $v_{\theta}$ which can be used to solve the ODE numerically. This results in the approximation $\log{\hat{p}_{\theta}}^{\mathrm{ODE}}\approx\log{p}$. The authors of [39] thus proposed to use $\log{\hat{p}_{\theta}}^{\mathrm{ODE}}$ to evaluate the objective (3.1.2). As the numerical results of [39] showed, the resulting VI approach has a high computational cost, but generates very accurate approximations of the true posterior. A follow-up paper [38] resolves the costly evaluation of $\log{\hat{p}_{\theta}}^{\mathrm{ODE}}$ by instead employing a surrogate, that is reminiscent of the evidence lower bound (ELBO) and was first derived in [97, Thm. 3]. The ELBO surrogate as used in [38, p. 5] is defined as follows,
$b_{\theta}^{\text{SDE}}(x_{0})=\mathbb{E}_{x\sim p_{T}(x|x_{0})}\left[\log\pi(x)\right]-\frac{1}{2}\int_{0}^{1}g(t)^{2}h(t,x_{0})\mathrm{d}t$ (3.12)
where
$h(t,x_{0})=\mathbb{E}_{x\sim p_{t}(x|x_{0})}\left[\|s_{\theta}(x,t)-\nabla_{x}\log p(x|x_{0})\|_{2}^{2}-\|\nabla_{x}\log p_{t}(x|x_{0})\|_{2}^{2}-\frac{2}{g(t)^{2}}\nabla\cdot f(x,t)\right].$ (3.13)
Under assumptions specified in [97, App. A], the authors show the following inequality [97, Thm. 3],
$b_{\theta}^{\text{SDE}}(x)\leq\log\hat{p}_{\theta}(x)\hskip 18.49988pt\forall x\in\mathcal{X}.$ (3.14)
In practice, $b_{\theta}^{\mathrm{SDE}}$ is approximated by replacing the integral (3.12) and the expectation inside the integrand term $h$ (3.13) with empirical Monte Carlo estimators using a finite number of samples, here reproduced from [38, Sec. 4.2]:
$b_{\theta}^{\mathrm{SDE}}\approx\tilde{b}_{\theta}^{\mathrm{SDE}}(x)$= $\frac{1}{N_{z}}\sum_{j=1}^{N_{z}}\log\pi\left(x_{j}^{\prime}\right)-\frac{1}{2N_{t}N_{z}}\sum_{i=1}^{N_{t}}Z\beta(t)^{2}\sum_{j=1}^{N_{z}}\Bigg[$ (3.15)
$\qquad\left\|\mathbf{s}_{\theta}(x_{ij}^{\prime},t_{i})+\frac{z_{ij}}{\beta(t_{i})}\right\|_{2}^{2}-\left\|\frac{z_{ij}}{\beta(t_{i})}\right\|_{2}^{2}-\frac{2}{g(t_{i})^{2}}\nabla_{x_{ij}^{\prime}}\cdot f(x_{ij}^{\prime},t_{i})\Bigg]$
s.t. $\qquad t_{i}\sim p(t),z_{ij}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),x_{ij}^{\prime}=\alpha(t_{i})x+\beta(t_{i})z_{ij}$
$x_{j}^{\prime}\sim\mathcal{N}(\alpha(T)x,\beta(T)^{2}\mathbf{I})\qquad\forall i=1,\ldots,N_{t},j=1,\ldots,N_{z},$
where we set $N_{t}=1,N_{z}=1$ and also use the re-weighted proposal distribution (c.f. [98, 38]) $p(t):=\frac{g(t)^{2}}{\beta(t)^{2}Z}$, where $Z=\int_{0}^{T}g(t)^{2}/\beta(t)^{2}\mathrm{d}t$. After this modification to (3.1.2), one instead solves
@@
$\tilde{\phi}\coloneqq\operatorname*{arg\,min}_{\phi}\left\{\mathbb{E}_{x\sim q_{\phi}}\left[-\log p(y\,|\,x)-\tilde{b}_{\theta}^{\mathrm{SDE}}(x)\right]-\mathcal{H}(q_{\phi})\right\},$ (3.16)
@@
where the ELBO term $\tilde{b}_{\theta}^{\mathrm{SDE}}$ is the Monte Carlo estimator (3.15). Variational inference naturally allows the parallel evaluation of the likelihood and the prior on image minibatches, consisting of multiple independent image samples, due to the ability to sample from the modeled distribution $q_{\phi}$. We further approximate the expectation in (3.16) using a batch of $B=4$ independent image samples $x_{1},\ldots,x_{B}\sim q_{\phi}$, used to evaluate both the likelihood term and $\tilde{b}_{\theta}^{\mathrm{SDE}}$. Heuristically, the ELBO term induces a surrogate score-based prior (SSP) with density ${\hat{p}_{\theta}}^{\mathrm{surr}}\propto\exp(b_{\theta}^{\mathrm{SDE}})$. The reduced computational time of (3.16) allows to lift the problem dimension to realistic imaging sizes.
Remark 3.1.
In general, one cannot expect the bound in 3.14 to be tight and an analytic characterization of the gap is an open problem. Empirically, it was shown that this surrogate prior is very effective, in the sense that for small dimensional examples $\tilde{\phi}\approx\phi^{\ast}$, i.e., the fitted distribution shows very good agreement with the variational approximation to the posterior [38, sec. 5].
In the case where the true posterior is Gaussian, [38, Fig. 3] shows that solutions to 3.16 coincide closely with solutions to the original problem 3.1.2. While [38, Fig. 4] highlights differences between using $\log\hat{p}_{\theta}$ and $\tilde{b}_{\theta}^{\text{SDE}}$ when applied for more complicated posterior distributions, the experiments still validate that the approach leads to a good posterior approximation.
@@
We now describe how this method can be extended to the case of blind inverse problems (3.4), using a standard variational Bayes approach for the joint posterior of image $x$ and latent parameter $\mathbf{r}$. While we are specifically interested in solving the position-blind ptychography problem (2.2), the method we derive can be used for general blind imaging inverse problems.
@@ 3.2.1 Generalized VI Approach for Blind Problems
Using (3.4), we generalize the VI approach to the joint posterior $p(x,\mathbf{r}\,|\,y)$ in the blind case. Under the non-restrictive assumption of independent priors $\log p(x,\mathbf{r})=\log{\hat{p}_{\theta}}(x)+\log p(\mathbf{r})$, the derivation of the VI objective (3.1.2) easily generalizes to
@@
Note that the variational distribution $q_{\phi}$ is now the joint distribution of image and parameters. Depending on the forward model and dimensionality, the joint optimization problem can exhibit strong non-convexity and be computationally demanding. Empirically, in preliminary experiments with optimizing $\chi,\rho$ jointly according to (3.17), we did not achieve any successful reconstructions. We therefore split the joint problem into sub-problems as follows. The problem can be provided with more structure by a standard mean-field assumption on the variational Bayesian posterior
@@
where $\phi=(\chi,\rho)$, with $\chi$ encoding the image posterior and $\rho$ the latent parameter posterior. Note that this comes at the cost of less accurately representing the correlation of uncertainties in image and parameters in $q_{\phi}$, but is necessary to make the optimization feasible at scale. Doing so also allows for higher flexibility; for instance, it lets us decouple the number of sequential optimization steps on $\rho$ and on $\chi$, a fact which we will practically make use of. We note that a detailed analysis of this approximation is an important open question for future research, but is outside of the scope of this work. By inspecting the optimality conditions of $\min_{\phi}L(\phi)=\min_{\chi,\rho}L(\chi,\rho)$ with respect to $\chi$ and $\rho$ separately, under the mean-field assumption, the posterior marginals obey the optimality conditions [13]
@@
$\rho_{k+1}$ $=\operatorname*{arg\,min}_{\rho\normalcolor}\mathcal{L}(\chi_{k+1},\rho)=\operatorname*{arg\,min}_{\rho}\left\{\mathbb{E}_{x\sim q_{\chi_{k+1}}}\mathbb{E}_{\mathbf{r}\sim q_{\rho}}\left[-\log p(y\,|\,x,\mathbf{r})-\log p(\mathbf{r})\right]-\mathcal{H}(q_{\rho})\right\}.$ (3.19b)
@@ 3.2.2 Blind RED-Diff
The RED-Diff method has been modified for tackling blind imaging inverse problems on the example of MRI with unknown off-resonance field map [1]. The authors there employed the same mean-field assumption (3.18) in order to separate optimization steps for the reconstructed image and parameters. We recall the framework established in [1] in the following, written in the notation we use here. As before, the objective for the image posterior can be rewritten using [77, Prop. 1] to arrive at a loss similar to (3.10), but with the likelihood conditioned on the current parameter estimate, i.e., (3.19a) becomes
@@
with a suitable weight $\tilde{\omega}(t)$; see [77, Prop. 2] for details. For the parameter update step (3.19b), the authors of [1] showed that assuming $p(\mathbf{r})$ and $q_{\rho}(\mathbf{r})$ are both Laplace distributions allows to obtain closed form representations of the relevant terms $p(\mathbf{r})$ and $\mathcal{H}(q_{\rho})$. Upon replacing $q_{\rho}$ with a point mass on a single $\mathbf{r}$, the entropy is, however, dropped again in [1] and the parameter update becomes
@@
We outline the resulting algorithm with our notation in Algorithm 2.
@@ 3.2.3 Blind surrogate score-based prior (SSP) method
$\chi_{k+1}=\operatorname*{arg\,min}_{\chi}\mathcal{L}(\chi,\rho_{k})=\operatorname*{arg\,min}_{\chi}\left\{\mathbb{E}_{x\sim q_{\chi}}\mathbb{E}_{\mathbf{r}\sim q_{\rho_{k}}}\left[-\log p(y\,|\,x,\mathbf{r})(x)-b_{\theta}^{\mathrm{SDE}}(x)\right]-\mathcal{H}(q_{\chi})\right\}.\normalcolor$ (3.23)
For medium- to large-scale imaging problems, $q_{\chi}$ can be a Gaussian distribution with diagonal covariance – more complex models were computationally prohibitive in our setup; see also [38] for scaling experiments. We solve (3.19) using inner loops of stochastic gradient descent, where the partial derivatives $\nabla_{\chi}\mathcal{L}$ and $\nabla_{\rho}\mathcal{L}$ are approximated using Monte Carlo estimators of the expectations and the typical reparametrization trick [64] whenever $q_{\chi},q_{\rho}$ are Gaussian. This sampling-based approach admits the use of a batch size $B\geq 1$ for each gradient evaluation, which can reduce the variance of estimated gradients. The surrogate prior term $b_{\theta}^{\mathrm{SDE}}(x)$ is numerically approximated with a Monte-Carlo estimator using a single forward-pass through $s_{\theta}$; see Section 3.1.4 and [38, sec. 4.2]. This leads to the variational Bayes approach summarized in Algorithm 1. Note that while the algorithm is presented generically, for simplicity, in practice we do not employ a particular stopping criterion but instead use a fixed number of iterations $N=l_{\mathrm{max}}$. For a clearer presentation, we formulate the updates as single gradient descent steps. However, in the practical implementation, lines 5 and 10 in Algorithm 1 are replaced by the Adam optimization scheme [65].